Compute the probabilities using the standard normal curve P(-1.89 < z < 1.68) 




1
Expert's answer
2022-02-14T15:53:53-0500

Solution:

P(-1.89 < z < 1.68)

= P(z < 1.68) - P(z < -1.89)

= P(z < 1.68) - P(z > 1.89)

= P(z < 1.68) - [1 - P(z "\le" 1.89)]

= 0.95352 - [1 - 0.97062] [Using z-score table]

= 0.92414